# On circumcentered direct methods for monotone variational inequality problems

### v1

@@ 4.1 Analysis of the direct method
Define $\bar{x}=\lim_{k\rightarrow\infty}P_{{\rm SOL}(F,C)}(\tilde{y}^{k})$. Then, either ${\rm SOL}(F,C)\neq\emptyset$ and $\{x^{k}\}$ converges to $\bar{x}$, or ${\rm SOL}(F,C)=\emptyset$ and $\lim_{k\rightarrow\infty}\|x^{k}\|=\infty$.
@@
Assume that ${\rm SOL}(F,C)\neq\emptyset$ and define $u^{k}=P_{{\rm SOL}(F,C)}(\tilde{y}^{k})$. Note that $u^{k}$ exists because ${\rm SOL}(F,C)$ is nonempty, closed, and convex by Proposition 2.13. By Lemma 4.2, $\{\tilde{y}^{k}\}$ is quasi-Fejér convergent to ${\rm SOL}(F,C)$ since ${\rm SOL}(F,C)\subset C$. Therefore, it follows from Proposition 2.16(ii) that $\{P_{{\rm SOL}(F,C)}(\tilde{y}^{k})\}$ is convergent. Let
$\bar{x}=\lim_{k\rightarrow\infty}P_{{\rm SOL}(F,C)}(\tilde{y}^{k})=\lim_{k\rightarrow\infty}u^{k}.$ (4.30)
By Lemma 4.2 and Theorem 4.4, $\{x^{k}\}$ is bounded and each of its cluster points belongs to ${\rm SOL}(F,C)$. Suppose that $\hat{x}\in{\rm SOL}(F,C)$ is a cluster point of $\{x^{k}\}$ and $\{x^{i_{k}}\}\rightarrow\hat{x}$. It suffices to show that $\bar{x}=\hat{x}$.
@@
$\langle\hat{x}-\bar{x},\tilde{y}^{j}-u^{j}\rangle\leq\langle u^{j}-\bar{x},\tilde{y}^{j}-u^{j}\rangle\leq\epsilon\|u^{j}-\bar{x}\|,$ (4.31)
for all $j$. Multiplying (4.31) by $\frac{\beta_{j}}{\eta_{j}\sigma_{k-1}}$ and summing from 00 to $k-1$, we get from Lemma 4.3(ii)
$\left\langle\hat{x}-\bar{x},x^{k}-\frac{1}{\sigma_{k-1}}\sum_{j=0}^{k-1}\frac{\beta_{j}}{\eta_{j}}u^{j}\right\rangle\leq\frac{\epsilon}{\sigma_{k-1}}\sum_{j=0}^{k-1}\frac{\beta_{j}}{\eta_{j}}\|u^{j}-\bar{x}\|.$ (4.32)
Define $\zeta_{k,j}:=\frac{1}{\sigma_{k}}\frac{\beta_{j}}{\eta_{j}}(k\geq 0,~{}0\leq j\leq k)$. In view of (4.27), $\lim_{k\rightarrow\infty}\zeta_{k,j}=0$ for all $j$. By (4.30),
$\sum_{j=0}^{k}\zeta_{k,j}=1,$ (4.33)
@@
$w^{k}=\sum_{j=0}^{k}\zeta_{k,j}u^{j}=\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}u^{j},$ (4.34)
@@
$\lim_{k\rightarrow\infty}\frac{1}{\sigma_{k-1}}\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}u^{j}=\bar{x},$ (4.35)
@@
$\lim_{k\rightarrow\infty}\frac{1}{\sigma_{k-1}}\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}\|u^{j}-\bar{x}\|=0,$ (4.36)
@@
If ${\rm SOL}(F,C)$ is empty then by Theorem 4.4 no subsequence of $\{x^{k}\}$ can be bounded, and hence $\lim_{k\rightarrow\infty}\|x^{k}\|=\infty$. ∎
@@
We remark that the sequence $\{x^{k}\}$ is a weighted average of the $k$ first elements of the sequence $\{\tilde{y}^{k}\}$. It is easy to check that if $\{\tilde{y}^{k}\}$ converges then so does $\{x^{k}\}$, but the converse statement does not necessarily hold. Indeed, the basic sequence in this algorithms is $\{\tilde{y}^{k}\}$, and the sequence $\{x^{k}\}$ was introduced only because we were not able to prove directly the convergence of $\{\tilde{y}^{k}\}$ ($\{x^{k}\}$ can be considered as an ergodic sequence related to $\{\tilde{y}^{k}\}$). In all our numerical experiments, the sequence $\{\tilde{y}^{k}\}$ did indeed converge, and obviously, faster than $\{x^{k}\}$, whose convergence is slowed down by the effect of the initial elements of $\{\tilde{y}^{k}\}$ in the computation of $x^{k}$. Hence, in Section 5 we report the results related to $\{\tilde{y}^{k}\}$ instead of $\{x^{k}\}$.
5 Numerical experiments
In this section, we present numerical experiments which evaluate the performance of our proposed Algorithms 1 and 2. We compare them with Algorithm 2 in [17], Algorithm 1 in [18], and the classical Extragradient method [45] for solving the variational inequality problem ${\rm VIP}(F,C)$. For clarity, we refer to these algorithms as Alg1, Alg2, Alg3, Alg4, and Extra, respectively. We recall that Alg1 is a circumcenter-based method intended to accelerate Alg3, and Alg2 acts in a similar way with respect to Alg4. Algorithms Alg1–Alg4 employ approximate projections onto convex sets, while Extra uses exact projections.
We consider the feasible set $C$ defined as the intersections of ellipsoids:
@@ 5 Numerical experiments
$C=\bigcap_{i=1}^{m}C_{i}.\quad C_{i}=\left\{x\in\mathds{R}^{n}:g_{i}(x)\leq 0\right\}=\left\{x\in\mathds{R}^{n}:x^{\top}A_{i}x+2(b^{i})^{\top}x-\alpha\leq 0\right\},$ (5.1)
@@
where each $A_{i}\in\mathds{R}^{n\times n}$ is a positive definite matrix, $b^{i}\in\mathds{R}^{n}$, and $\alpha\in\mathds{R}$. These ellipsoids are generated following the approach in [10, Section 5], thus ensuring the existence of a Slater point, required for the inner loop of both Alg2 and Alg4.
@@
For computing the circumcenter in (1.23), we use the gradient of $g_{i}$, which is given by $\nabla g_{i}(x)=2A_{i}x+b^{i}$. However, computing the exact projection onto $\bigcap_{i=1}^{m}C_{i}$ required in the Extragradient steps (1.6) and (1.7) is nontrivial. We employ Dykstra’s algorithm [29] and ADMM [41, Algorithm 6] for computing the projection onto each $C_{i}$, which turn out to work well for this purpose.
Since Alg2 [17] and Alg4 [18] require halfspaces separating $C$ from the current iterate, we define the function
$\hat{g}(x):=\max_{i\in\{1,\ldots,m\}}g_{i}(x),$ (5.2)
and construct the halfspace $\hat{H}(x)$, defined as
$\hat{H}(x):=\left\{z\in\mathds{R}^{n}:\hat{g}(z)\geq\hat{g}(x)+\langle u,z-x\rangle\right\},$ (5.3)
with $u\in\partial\hat{g}(x)$. It is well known that $u$ can be taken as $\nabla g_{k}(x)$, where $k$ is any index which realizes the maximum in (5.2). We mention that $\hat{H}(x)$ separates the point $x$ from the feasible set $C$.
We perform two kinds of experiments:
  • Approximate Projection Comparison: We compare Alg1–Alg4 in terms of CPU time and iteration counts, for highlighting the acceleration effect of the circumcenter approach in Alg1 and Alg2, with respect to the other two algorithms.
  • Exact vs. Approximate Projection: We compare Alg1, Alg2, and Extra in terms of CPU time and final error, in order to investigate the practical impact of using exact versus approximate projections.
The computational experiments were performed on an Intel Core i7-12700H 2.30GHz with 32 GB of RAM running Windows 10 Professional version using Matlab 2015b, and are fully available at
https://github.com/Yanziszj/CRM-VIP. The following conditions were used:
  1. 1.
    The maximum of iterations for all the algorithms is 30000.
  2. 2.
    The stopping criteria for Alg1 and Alg3 are
    $\|x^{k+1}-x^{k}\|\leq\epsilon~{}{\rm or}~{}k>30000.$ (5.4)
    The stopping criteria for Alg2 and Alg4 are
    $\|z^{k+1}-y^{k}\|\leq\epsilon~{}{\rm or}~{}k>30000.$ (5.5)
    The stopping criteria for Extra are
    $\|x^{k}-y^{k}\|\leq\epsilon~{}{\rm or}~{}k>30000,$ (5.6)
    where $\epsilon=10^{-6}$ and $k$ is the iteration index.
  3. 3.
    The parameters in ADMM are the same as in [41, Algorithm 6]. The stopping criterion of Dykstra’s Algorithm is the same as in [19, Sect. 4]
  4. 4.
    For determinig whether algorithms using approximate projection properly approach the solution set of ${\rm VIP}(F,C)$, we compute the error at iteration $k$ of each algorithm in the second part of our experiments by the following rule:
    ${\rm error}=\|x^{k}-P_{C}(x^{k}-\alpha F(x^{k}))\|,$ (5.7)
    with $\alpha>0$. We choose as $\alpha=0.1$. It is well known that if $x\in{\rm SOL}(F,C)$, then $x$ is also a fixed point of $\left[P_{C}\circ(I-\alpha F)\right](\cdot)$ for any positive $\alpha$.
  5. 5.
    For each pair $(n,m)$, for $n\in\{5,10,20\}$ and $m\in\{2,5,10\}$ we generate 20 examples and run the test for each example 10 times.
  6. 6.
    We choose $\beta_{k}=\frac{1}{k}$ for Alg1-Alg4 according to (3.4) and (3.5) in Algorithm 1, (4.1) and (4.2) in Algorithm and similar assumptions in [17] and [18]. We take a constant stepsize $\alpha=0.05$ for Extra.
5.1 Examples of operators in ${\rm VIP}(F,C)$
In order to test Alg1 and Ag2 for ${\rm VIP}(F,C)$ with paramonotone and monotone operators, we consider the following three families of operators.
Example 1: First, we test examples with operators $F$ which are gradients of convex functions. In these cases, ${\rm VIP}(F,C)$ is equivalent to the problem of minimizing the convex function with gradient $F$ over $C$. Define
$F(x)=Ax+G(x)+c$ (5.8)
@@ 5.1 Examples of operators in ${\rm VIP}(F,C)$
where $G(x)_{i}=b_{i}x_{i}^{3}$ with $b,c\in\mathds{R}^{n},b\geq 0,c\neq 0$. We generate randomly $A\in\mathds{R}^{n\times n}$ as a symmetric, positive semidefinite matrix. Note that $F(x)$ is the gradient of the convex function $f(x)=\frac{1}{2}x^{t}Ax+cx+\sum_{i=1}^{n}b_{i}x_{i}^{4}$.
@@
Example 2: It is known that the gradient of a convex function is a paramonotone operator [23] but the inverse is not true. We consider now operators which are not gradients of convex functions. These operators ae defined as:
$F(x)=Ax+c,$ (5.9)
where the matrix $A\subset\mathds{R}^{n\times n}$ is defined as
$A=\left[\begin{array}[]{cc}A_{1}&0\\ 0&A_{2}\end{array}\right].$ (5.10)
with a symmetric positive semidefinite matrix $A_{1}\in\mathds{R}^{n_{1}\times n_{1}}$ and a nonsymmetric positive semidefinite matrix $A_{2}\in\mathds{R}^{n_{2}\times n_{2}}$. $F$ is not a gradient, because $A$ is the Jacobian matrix of $F$ and Jacobian matrices of gradients are symmetric.

### v2

@@ 4.1. Analysis of the direct method
If $F$ and $C$ satisfy Assumptions 1.8 and 1.9, (i.e. $F$ is continuous and monotone), then, either $\operatorname{SOL}(F,C)\neq\emptyset$ and $\{x^{k}\}$ converges to $\bar{x}$, where $\bar{x}=\lim_{k\rightarrow\infty}P_{\operatorname{SOL}(F,C)}(\tilde{y}^{k})$ or $\operatorname{SOL}(F,C)=\emptyset$ and $\lim_{k\rightarrow\infty}\|x^{k}\|=\infty$.
@@
Assume that $\operatorname{SOL}(F,C)\neq\emptyset$ and define $u^{k}=P_{\operatorname{SOL}(F,C)}(\tilde{y}^{k})$. Note that $u^{k}$ exists because $\operatorname{SOL}(F,C)$ is nonempty, closed, and convex by Proposition 2.13. By Lemma 4.3, $\{\tilde{y}^{k}\}$ is quasi-Fejér convergent to $\operatorname{SOL}(F,C)$ since $\operatorname{SOL}(F,C)\subset C$. Therefore, it follows from Proposition 2.16(ii) that $\{P_{\operatorname{SOL}(F,C)}(\tilde{y}^{k})\}$ is convergent. Let
(4.28) $\bar{x}=\lim_{k\rightarrow\infty}P_{\operatorname{SOL}(F,C)}(\tilde{y}^{k})=\lim_{k\rightarrow\infty}u^{k}.$
By Lemma 4.3 and Lemma 4.5, $\{x^{k}\}$ is bounded and each of its cluster points belongs to $\operatorname{SOL}(F,C)$. Suppose that $\hat{x}\in\operatorname{SOL}(F,C)$ is a cluster point of $\{x^{k}\}$ and $\{x^{i_{k}}\}\rightarrow\hat{x}$. It suffices to show that $\bar{x}=\hat{x}$.
@@
(4.29) $\langle\hat{x}-\bar{x},\tilde{y}^{j}-u^{j}\rangle\leq\langle u^{j}-\bar{x},\tilde{y}^{j}-u^{j}\rangle\leq\epsilon\|u^{j}-\bar{x}\|,$
for all $j$. Multiplying (4.29) by $\frac{\beta_{j}}{\eta_{j}\sigma_{k-1}}$ and summing from 0 to k1k-1, we get from Lemma 4.4(ii)
(4.30) $\left\langle\hat{x}-\bar{x},x^{k}-\frac{1}{\sigma_{k-1}}\sum_{j=0}^{k-1}\frac{\beta_{j}}{\eta_{j}}u^{j}\right\rangle\leq\frac{\epsilon}{\sigma_{k-1}}\sum_{j=0}^{k-1}\frac{\beta_{j}}{\eta_{j}}\|u^{j}-\bar{x}\|.$
Define $\zeta_{k,j}\coloneqq\frac{1}{\sigma_{k}}\frac{\beta_{j}}{\eta_{j}}(k\geq 0,~0\leq j\leq k)$. In view of (4.25), $\lim_{k\rightarrow\infty}\zeta_{k,j}=0$ for all $j$. By (4.28),
(4.31) $\sum_{j=0}^{k}\zeta_{k,j}=1,$
@@
(4.32) $w^{k}=\sum_{j=0}^{k}\zeta_{k,j}u^{j}=\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}u^{j},$
@@
(4.33) $\lim_{k\rightarrow\infty}\frac{1}{\sigma_{k-1}}\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}u^{j}=\bar{x},$
@@
(4.34) $\lim_{k\rightarrow\infty}\frac{1}{\sigma_{k-1}}\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}\|u^{j}-\bar{x}\|=0,$
@@
If $\operatorname{SOL}(F,C)$ is empty then by Lemma 4.5 no subsequence of $\{x^{k}\}$ can be bounded, and hence $\lim_{k\rightarrow\infty}\|x^{k}\|=\infty$. ∎
@@
We remark that the sequence $\{x^{k}\}$ is a weighted average of the $k$ first elements of the sequence $\{\tilde{y}^{k}\}$. It is easy to check that if $\{\tilde{y}^{k}\}$ converges then so does $\{x^{k}\}$, but the converse statement does not hold. Indeed, if we consider $\operatorname{VIP}(F,C)$ in $\mathds{R}^{2}$, taking $F$ as the $\pi/2$-rotation defined in Subsection 1.2 and $C$ as the unit disk, then the only solution of $\operatorname{VIP}(F,C)$ is 0, but the sequence $\{\tilde{y}^{k}\}$ does not converge; rather, it accumulates on the boundary of $C$. The averaged sequence $\{x^{k}\}$, on the other hand, does converge to 0. In all our numerical experiments in Section 5, we monitor both $\{\tilde{y}^{k}\}$ and $\{x^{k}\}$. Whenever $\{\tilde{y}^{k}\}$ converges (which happens almost always), we report the results related to $\{\tilde{y}^{k}\}$ instead of $\{x^{k}\}$. Obviously, in these cases $\{\tilde{y}^{k}\}$ converges faster than $\{x^{k}\}$, whose convergence is slowed down by the effect of the initial elements of $\{\tilde{y}^{k}\}$ in the computation of $x^{k}$.
5. Numerical experiments
In this section, we present numerical experiments to evaluate the performance of our proposed Algorithms 1 and 2, which we refer to as CRM-VIP1 and CRM-VIP2, respectively.
5.1. Setup and methodology
We conduct experiments comparing our proposed algorithms with established methods for solving variational inequality problems (VIPs). To put our methods to the test, beyond Algorithm 2 in [17] (referred to as BI1), Algorithm A in [18] (denoted by BI2), we selected the classical Extragradient method [46] (called here EGM), and Algorithm 4.2 of [49] (abbreviated as Mal-Adap). We recall that CRM-VIP1 and CRM-VIP2 are circumcenter-based methods intended to accelerate BI1 and BI2, respectively. The methods CRM-VIP1, CRM-VIP2, BI1, and BI2 employ approximate projections (based on separating halfspaces), while EGM and Mal-Adap use exact projections.
We consider the feasible set $C$ defined as the intersection of ellipsoids:
@@ 5.1. Setup and methodology
(5.1) $C=\bigcap_{i=1}^{m}C_{i},\quad C_{i}\coloneqq\left\{x\in\mathds{R}^{n}:g_{i}(x)\leq 0\right\},\text{ where }g_{i}(x)=\left\langle x,A_{i}x\right\rangle+2\left\langle b^{i},x\right\rangle-\alpha,$
@@
where each $A_{i}\in\mathds{R}^{n\times n}$ is a positive definite matrix, $b^{i}\in\mathds{R}^{n}$, and $\alpha\in\mathds{R}$. These ellipsoids are generated following the approach in [11]*Section 5, ensuring the existence of a Slater point, which is required for the inner loop of both BI2 and CRM-VIP2.
@@
For computing the circumcenter step in (1.20), we use the gradient of $g_{i}$, given by $\nabla g_{i}(x)=2A_{i}x+2b^{i}$. However, computing the exact projection onto $C=\bigcap_{i=1}^{m}C_{i}$, as required by EGM and Mal-Adap, is nontrivial. We employ Dykstra’s algorithm [31] for computing the projection onto $C$, where each subprojection onto $C_{i}$ is computed via ADMM as described in [43]*Algorithm 6.
Importantly, CRM-VIP1 and CRM-VIP2 share the same codebase as BI1 and BI2, respectively, differing only in the projection step: CRM-VIP1 and CRM-VIP2 replace the approximate projection with the circumcenter acceleration. This design ensures a fair comparison, isolating the effect of the circumcenter acceleration.
Given $x\notin C$, we define
(5.2) $\hat{g}(x):=\max_{i\in\{1,\ldots,m\}}g_{i}(x),$
and let $k$ be any index attaining the maximum. The halfspace
(5.3) $\hat{H}(x):=\left\{z\in\mathds{R}^{n}:g_{k}(x)+\langle\nabla g_{k}(x),z-x\rangle\leq 0\right\},$
separates $x$ from $C$, and the projection onto $\hat{H}(x)$ admits a closed form. The existence of a Slater point, guaranteed by the ellipsoid generation procedure, is required for the inner loops of CRM-VIP2 and BI2.
The experiments are organized into three scenarios according to problem size:
  1. a.
    Scenario A (Small): $n\in\{5,10\}$ and $m\in\{2,5\}$. All six methods are tested.
  2. b.
    Scenario B (Medium): $n\in\{50,100\}$ and $m\in\{5,8\}$. Only approximate projection methods (CRM-VIP1, CRM-VIP2, BI1, BI2) are tested.
  3. c.
    Scenario C (Large): $n\in\{100,200,500\}$ and $m\in\{20,30,50\}$. Only approximate projection methods are tested.
This structure allows us to assess both the accuracy trade-off between exact and approximate projections (Scenario A) and the scalability of approximate projection methods as well as checking the benefits of circumcenter-type methods (Scenarios B and C).
All experiments were implemented in Julia 1.12 [19] and executed on a MacBook Air M4 with 24 GB RAM and 10-core CPU. The source code is available at https://github.com/lrsantos11/CRM_VIP.jl. The following conditions were used:
  1. (i)
    The maximum number of iterations is 100 000100\,000 for Scenario A and 300 000300\,000 for Scenarios B and C.
  2. (ii)
    The stopping criteria vary by method:
    • For CRM-VIP1 and BI1:
      (5.4) $\frac{\|x^{k+1}-x^{k}\|}{\max\{\|x^{k}\|,1\}}\leq\varepsilon.$
    • For CRM-VIP2 and BI2, the algorithm stops when either the above criterion holds for the ergodic sequence $\{x^{k}\}$, or
      (5.5) $\|z^{k}-\tilde{y}^{k}\|\leq\varepsilon.$
    • For EGM:
      (5.6) $\|x^{k}-y^{k}\|\leq\varepsilon.$
    • For Mal-Adap, we use the residual function from [49]:
      (5.7) $r(x^{k},y^{k})=\|y^{k}-P_{C}(y^{k}-\lambda F(y^{k}))\|+\|x^{k}-y^{k}\|\leq\varepsilon.$
    We set $\varepsilon=10^{-6}$ for Scenarios A and B, and $\varepsilon=10^{-5}$ for Scenario C.
  3. (iii)
    We set $\beta_{k}=1/k^{0.9}$ for CRM-VIP1, CRM-VIP2, BI1, and BI2, which satisfies (1.22)–(1.23) and the analogous conditions in [17, 18]. For Mal-Adap, the step size is adapted at each iteration as described in [49]*Alg. 4.2, with initial step size $\lambda_{0}=0.5/L$, where $L$ is an estimate of the Lipschitz constant of $F$.
  4. (iv)
    For each configuration $(n,m)$, we generate $10$ random instances for Scenarios A and B, and $5$ instances for Scenario C.
  5. (v)
    In all scenarios we run the experiments in $10$ threads to parallelize the execution of different instances. When the instance had more than $8$ sets, the computation of the circumcenter step was also parallelized Julia package ThreadsX.jl.
  6. (vi)
    Wall-clock times are measured using the BenchmarkTools.jl package [25], which reports the minimum time over multiple executions, providing reliable and reproducible timing measurements.
5.2. Test problems
We evaluate the algorithms on three families of operators $F$ for $\operatorname{VIP}(F,C)$, designed to test different theoretical properties: gradient operators (paramonotone), non-gradient paramonotone operators, and monotone but non-paramonotone operators.
Example 5.1 (Gradient of convex function – paramonotone).
We consider operators $F$ that are gradients of convex functions. In this case, solving $\operatorname{VIP}(F,C)$ is equivalent to minimizing a convex function over $C$. The operator is defined as
(5.8) $F(x)=Ax+G(x)+c,$
@@ 5.2. Test problems
where $G(x)_{i}=b_{i}x_{i}^{3}$ with $b\in\mathds{R}^{n}$, $b\geq 0$, and $c\in\mathds{R}^{n}$, $c\neq 0$. The matrix $A\in\mathds{R}^{n\times n}$ is generated randomly as symmetric positive semidefinite (via $A=MM^{\top}$ for a random matrix $M$). Note that $F$ is the gradient of the convex function
$f(x)=\frac{1}{2}\left\langle x,Ax\right\rangle+\left\langle c,x\right\rangle+\frac{1}{4}\sum_{i=1}^{n}b_{i}x_{i}^{4}.$
Since gradients of convex functions are paramonotone [24], all algorithms are expected to converge.
Example 5.2 (Paramonotone, non-gradient operator).
@@
We now consider operators that are paramonotone but not gradients of convex functions. The operator is defined as
(5.9) $F(x)=Ax+c,$
where $A\in\mathds{R}^{n\times n}$ has the block-diagonal structure
(5.10) $A=\begin{bmatrix}A_{1}&0\\ 0&A_{2}\end{bmatrix}$,
with $A_{1}\in\mathds{R}^{n_{1}\times n_{1}}$ symmetric positive semidefinite and $A_{2}\in\mathds{R}^{n_{2}\times n_{2}}$ nonsymmetric positive semidefinite (where $n_{1}+n_{2}=n$). We construct $A_{2}$ following [49] as $A_{2}=M_{2}M_{2}^{\top}+B+D$, where $B$ is skew-symmetric and $D$ is diagonal with positive entries. Since $A$ is not symmetric, $F$ is not a gradient. However, $F$ is paramonotone because $\operatorname{rank}(A+A^{\top})=\operatorname{rank}(A)$.
Example 5.3 (Monotone, non-paramonotone operator).